Using Haskell

Write a tripleDistance function that takes two 3-tuples and finds the cartesian distance

LaTeX: d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2 + (z_2 - z_1)^2}

use Haskell's sqrt, ^, +, *, - as needed
Answer #1

`Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

Note: Brother sometimes while uploading on HomeworkLib the indentations change. So, I request you to verify it with screenshot once.

tripleDistance(x1 , y1, z1) (x2 , y2, z2)= sqrt (x'*x' + y'*y'+ z'*z')
where
x' = x1 - x2
y' = y1 - y2
z' = z1 - z2
  
1 tripleDistance(x1 z1) (x2 z2)-sqrt (x.*x y*y+ z*z) 1 yl, y2, + , , where 2

Kindly revert for any queries

Thanks.